# Deep Optimal Transport for Domain Adaptation on SPD Manifolds

### v4

@@ Abstract
The machine learning community has shown increasing interest in addressing the domain adaptation problem on symmetric positive definite (SPD) manifolds. This interest is primarily driven by the complexities of neuroimaging data generated from brain signals, which often exhibit shifts in data distribution across recording sessions. These neuroimaging data, represented by signal covariance matrices, possess the mathematical properties of symmetry and positive definiteness. However, applying conventional domain adaptation methods is challenging because these mathematical properties can be disrupted when operating on covariance matrices. In this study, we introduce a novel geometric deep learning-based approach utilizing optimal transport on SPD manifolds to manage discrepancies in both marginal and conditional distributions between the source and target domains. We evaluate the effectiveness of this approach in three cross-session brain-computer interface scenarios and provide visualized results for further insights. The GitHub repository of this study can be accessed at https://github.com/GeometricBCI/Deep-Optimal-Transport-for-Domain-Adaptation-on-SPD-Manifolds.
Index Terms:
@@
Domain Adaptation, Optimal Transport, Geometric Deep Learning, Neural Signal Processing
I Introduction
In cross-domain problems, the distribution of source data often differs from that of target data, leading to what is widely recognized as a domain adaptation problem in the machine learning community. This issue has gained considerable attention over the past decade [1, 2, 3]. For example, signals in an online testing set frequently differ from those in the calibration set, despite being generated by the same process, highlighting a common domain adaptation challenge.
In domain adaptation, a domain $\mathcal{D}$ consists of a feature space $X$ and a marginal probability distribution $\mathbb{P}(X)$, denoted as $\mathcal{D}=\{X,\mathbb{P}(X)\}$. Meanwhile, a task $\mathcal{T}$ comprises a label space $Y$ and a conditional distribution $\mathbb{Q}(Y|X)$, represented by $\mathcal{T}=\{Y,\mathbb{Q}(Y|X)\}$. Domain adaptation aims to build a classifier in the target domain $\mathcal{D}_{T}$ by leveraging the information obtained from the source domain $\mathcal{D}_{S}$, where $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ and $\mathcal{T}_{S}\neq\mathcal{T}_{T}$. In particular, $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ yields either $X_{S}\neq X_{T}$ or $\mathbb{P}_{S}(X_{S})\neq\mathbb{P}_{T}(X_{T})$, whereas $\mathcal{T}_{S}\neq\mathcal{T}_{T}$ results in either $Y_{S}\neq Y_{T}$ or $\mathbb{Q}_{S}(Y_{S}|X_{S})$ or $\mathbb{Q}_{T}(Y_{T}|X_{T})$.
The domain adaptation problem can be broadly categorized into unsupervised and semi-supervised scenarios. In the unsupervised scenario, no labels are available in the target domain, while in the semi-supervised scenario, a limited number of labeled instances are available in the target domain to guide the adaptation process. For instance, the variability of electroencephalography (EEG) signals often results in unreliable and non-robust performance when transitioning from calibration to feedback phases, as these sessions exhibit distinct distributions. This variability in EEG-based classification exemplifies a typical domain adaptation problem [4]. In response to this issue, a covariate shift adaptation approach has been proposed in machine learning society [5]. This approach assumes that the difference between domains is characterized by a change in the feature space, while the conditional distributions remain unchanged.
TABLE I: A comparison of various OT-DA Frameworks. $X_{S}$ and $X_{T}$ are feature spaces of the source and target domains respectively. $\mathbb{P}$ and $\mathbb{Q}$ represent the marginal probability distribution and conditional probability distribution respectively.
Framework Cost function Underlying Space Transformation Domain Adaptation Scenario
OT-DA[6, 7] $c(x,\bar{x})=||x-\bar{x}||_{\ell_{2}}^{2}.$ Euclidean Affine $X_{S}\neq X_{T}$.
JDOT-DA[8, 9] $c(x,\bar{x})=||x-\bar{x}||_{\ell_{2}}^{2}$ Euclidean Affine $X_{S}\neq X_{T}$, $\mathbb{Q}_{S}(Y_{S}|X_{S})\neq\mathbb{Q}_{T}(Y_{T}|X_{T})$.
OT-DA on SPD manifolds[10] $c(x,\bar{x})=||x-\bar{x}||_{\ell_{2}}^{2}$ ($\mathcal{S}_{++}$, $g^{AIRM}$) Bi-Map $X_{S}\neq X_{T}$.
OT-DA on SPD manifolds (Proposed) $c(x,\bar{x})=d_{g^{LEM}}^{2}(x,\bar{x})$ ($\mathcal{S}_{++}$, $g^{LEM}$) Neural Networks $X_{S}=X_{T}$, $\mathbb{P}_{S}(X_{S})\neq\mathbb{P}_{T}(X_{T})$, and
$\mathbb{Q}_{S}(Y_{S}|X_{S})\neq\mathbb{Q}_{T}(Y_{T}|X_{T})$.
Among various methods that have been proposed to address the domain adaptation problem across numerous engineering scenarios, Optimal Transport (OT) has emerged as a novel and effective solution for aligning source and target domains [6, 7]. The OT approach introduces regularized optimal transport, allowing a classifier to learn from labeled source data and apply this knowledge to target data. The goal of the Optimal Transport-Domain Adaptation (OT-DA) framework is to minimize the overall transportation costs between two distributions, quantified by the Wasserstein metric. It seeks to find a push-forward transformation $\varphi:X_{S}\longmapsto X_{T}$ that satisfies $\mathbb{P}_{S}(X_{S})=\mathbb{P}_{T}(\varphi(X_{S}))$. The strengths of the OT-DA framework lie in its simplified evaluation of empirical data distribution and its enhanced ability to leverage the underlying geometry of discrete samples [6].
While broad in scope, the OT-DA framework can benefit from enhancements derived from insights into specific applications. For instance, in EEG-based motor imagery classification, EEG signals are processed as (average) spatial covariance matrices after segmentation in time and averaging over trials [11, 12]. This leads to the cross-session scenario of motor imagery classification taking place in the space of covariance matrices, which forms a symmetric and positive definite (SPD) manifold if endowed with a Riemannian metric. In this context, Yair et al. utilized squared $\ell_{2}$ distance as the cost function $c(x,\bar{x})=||x-\bar{x}||_{\ell_{2}}^{2}$ in the OT-DA framework, deriving a closed-form solution through Brenier’s polar factorization [10].
Although Yair’s work pioneered the OT-DA framework on SPD manifolds, it presents certain theoretical and practical shortcomings. Theoretically, their chosen cost function, based on Euclidean distance, undermines the consistency of mathematical formulation in the view of differential geometry. Practically, their methodology fails to account for disparities within conditional distributions. Additionally, the unique challenges inherent to the problem, such as the scarcity of labeled data during the feedback phase and the divergence of spatial covariance matrices from distinct frequency bands, pose significant modeling difficulties. These challenges necessitate the development of a more adaptable OT-DA framework to effectively address problems on SPD manifolds.
To address these challenges, we consider a more applicable scenario where $X_{S}=X_{T}$, $Y_{S}=Y_{T}$, $\mathbb{P}_{S}(X_{S})\neq\mathbb{P}_{T}(X_{T})$, and $\mathbb{Q}_{S}(Y_{S}|X_{S})\neq\mathbb{Q}_{T}(Y_{T}|X_{T})$. This situation is typically handled through joint distribution adaptation, which concurrently aligns both marginal and conditional distributions across domains, even when dealing with implicit probability distributions, without relying on labeled information from the target domain [13]. To facilitate modeling of $X_{S}=X_{T}$, we formulate the space of SPD matrices using the Log-Euclidean metric $g^{LEM}$ on SPD manifolds. The Log-Euclidean metric is a commonly used Riemannian metric on SPD manifolds, allowing operations to be performed in the logarithmic domain where the structure is Euclidean [14]. Moreover, in this context, parallel transport on SPD manifolds with the Log-Euclidean metric simplifies to an identity map, ensuring that the feature space of the source and target domains, which is the tangent space at the identity of SPD manifolds, are the same.
In this study, we propose a novel geometric deep learning architecture called Deep Optimal Transport (DOT) to address the joint distribution adaptation problem on SPD manifolds, utilizing a squared Log-Euclidean distance as the cost function in optimal transport, i.e., $c(x,\bar{x})=d_{g^{LEM}}^{2}(x,\bar{x})$. DOT is predicated on the assumption that information transfer between domains can be effectively estimated through optimal transport. The goal of DOT is to find a transformation on SPD manifolds $\varphi:X_{S},X_{T}\longmapsto Z\subset\mathcal{S}_{++}$ such that $\mathbb{P}_{S}(\varphi(X_{S}))=\mathbb{P}_{T}(\varphi(X_{T}))$ and $\mathbb{Q}_{S}(Y_{S}|\varphi(X_{S}))=\mathbb{Q}_{T}(Y_{T}|\varphi(X_{T}))$. To evaluate the effectiveness of DOT, we focus on cross-session scenarios in motor imagery classification. This problem is notably challenging and often inadequately addressed, primarily due to the extensive patterns of synchronized neuronal activity that continuously shift over time. This results in significant variability in brain signals captured by EEG devices across different sessions. In our evaluation, we calibrate the model in one session and then test it in another, encompassing both unsupervised and semi-supervised domain adaptation scenarios.
@@ I Introduction
The remainder of this paper is organized as follows: Section II delves into the Log-Euclidean geometry, the discrete OT problems, joint distribution optimal transport, and SPD matrix-valued motor imagery classifier. Section III presents the OT-DA framework on SPD manifolds and proposes the architecture of DOT. In Section IV, we will employ DOT on the motor imagery classification. Section V exhibits results on synthetic and real-world datasets. A comparison of various frameworks is summarized in Table I.
II PRELIMINARY
@@
A matrix $S\in\mathbb{R}^{n\times n}$ is symmetric and positive-definite if it is symmetric and all its eigenvalues are positive, i.e., $S=S^{T}$, and $x^{T}\cdot S\cdot x>0$, for all $x\in\mathbb{R}^{n}\setminus\{{\mathbf{0}}\}$. The Frobenius inner product and Frobenius norm on $m\times n$ matrices $A$ and $B$ are defined as $\langle\,A,B\,\rangle_{\mathcal{F}}:=\operatorname{Tr}{(A^{\top}\cdot B)}$ and $||\,A\,||_{\mathcal{F}}^{2}:=\langle\,A,A\,\rangle_{\mathcal{F}}$ respectively.
II-A Log-Euclidean Metric
@@
In this study, another Riemannian metric on SPD manifolds, affine-invariant Riemannian metric [16], is introduced for comparison with $g^{LEM}$. SPD manifold equipped with this metric is denoted as ($\mathcal{S}_{++}$, $g^{AIRM}$). Formally, for any $P$, it is defined as $g^{AIRM}_{P}(v,w):=\langle\,P^{-\frac{1}{2}}\,v\,P^{-\frac{1}{2}},P^{-\frac{1}{2}}\,w\,P^{-\frac{1}{2}}\,\rangle_{\mathcal{F}}$, where tangent vectors $v,w\in\mathcal{T}_{P}\mathcal{S}_{++}$. ($\mathcal{S}_{++}$, $g^{AIRM}$) also exhibit nice properties, such as affine invariance, scale invariance, and inversion invariance. Since these properties are not the primary focus of this study, please refer to details in [17].
II-B Discrete Optimal Transport
@@
Optimal Transport (OT) seeks the minimal cost transport plan between source and target distributions [18]. In this study, we consider the case where these distributions are available only through a finite number of samples. Given two measurable spaces $X$ and $\bar{X}$, an empirical distribution of the measure $\mu\in X$ and $\nu\in\bar{X}$ are written as $\mu:=\sum_{i\in I}p_{i}\delta_{x_{i}}$ and $\nu:=\sum_{j\in J}q_{i}\delta_{\bar{x}_{i}}$, where discrete variable $X$ takes values $\{x_{i}\}_{i\in I}\in X$, countable index set $I$ is with weights $\sum_{i\in I}p_{i}=1$. $\bar{X}$ takes values $\{\bar{x}_{j}\}_{j\in J}\in\bar{X}$, countable index set $J$ is with weights $\sum_{j\in J}q_{j}=1$. Dirac measure $\delta_{s}$ denotes a unite point mass at the point $s\in X\text{ or }\bar{X}$. The discrete Kantorovich formalization of the OT problem is given as $\min_{\gamma\in\prod(\mu,\nu)}\big{\langle}\gamma,c(x,\bar{x})\big{\rangle}_{\mathcal{F}}$, where transportation plan $\gamma\in\prod(\mu,\nu):=\{\gamma\in\mathbb{R}_{\geq 0}^{|X|\times|\bar{X}|}\big{|}\gamma\mathbbm{1}_{|X|}=\mu,\text{ and }\gamma^{\top}\mathbbm{1}_{|\bar{X}|}=\nu\}$, and $\mathbbm{1}_{d}$ is a $d-$dimensional all-one vector.
II-C Joint Distribution Optimal Transport

### v5

@@ Abstract
Recent progress in geometric deep learning has drawn increasing attention from the machine learning community toward domain adaptation on symmetric positive definite (SPD) manifolds—especially for neuroimaging data that often suffer from distribution shifts across sessions. These data, typically represented as covariance matrices of brain signals, inherently lie on SPD manifolds due to their symmetry and positive definiteness. However, conventional domain adaptation methods often overlook this geometric structure when applied directly to covariance matrices, which can result in suboptimal performance. To address this issue, we introduce a new geometric deep learning framework that combines optimal transport theory with the geometry of SPD manifolds. Our approach aligns data distributions while respecting the manifold structure, effectively reducing both marginal and conditional discrepancies. We validate our method on three cross-session brain-computer interface datasets—KU, BNCI2014001, and BNCI2015001—where it consistently outperforms baseline approaches while maintaining the intrinsic geometry of the data. We also provide quantitative results and visualizations to better illustrate the behavior of the learned embeddings.
keywords:
@@
Domain Adaptation , Optimal Transport , Geometric Deep Learning , Riemannian Manifolds
1 Introduction
Cross-domain problems arise when the distribution of source data diverges from that of target data, a challenge widely studied as domain adaptation in machine learning [1, 2, 3]. For instance, even if signals in an online testing set and a calibration set originate from the same generative process, their distributions often differ significantly. Formally, a domain $\mathcal{D}$ is defined by a feature space $X$ and its marginal probability distribution $\mathbb{P}(X)$, denoted as $\mathcal{D}=\{X,\mathbb{P}(X)\}$. A task $\mathcal{T}$ consists of a label space $Y$ and the conditional distribution $\mathbb{Q}(Y|X)$, expressed as $\mathcal{T}=\{Y,\mathbb{Q}(Y|X)\}$. Domain adaptation aims to transfer knowledge from a source domain $\mathcal{D}_{S}$ and task $\mathcal{T}_{S}$ to improve model performance on a target domain $\mathcal{D}_{T}$ and task $\mathcal{T}_{T}$, where $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ and $\mathcal{T}_{S}\neq\mathcal{T}_{T}$, specifically as follows:
  • $\mathcal{D}_{S}\neq\mathcal{D}_{T}$ implies either $X_{S}\neq X_{T}$ (e.g., differing feature dimensions) or $\mathbb{P}_{S}(X_{S})\neq\mathbb{P}_{T}(X_{T})$ (e.g., covariate shift).
  • $\mathcal{T}_{S}\neq\mathcal{T}_{T}$ indicates either $Y_{S}\neq Y_{T}$ (e.g., label space mismatch) or $\mathbb{Q}_{S}(Y_{S}|X_{S})\neq\mathbb{Q}_{T}(Y_{T}|X_{T})$ (e.g., label shift).
Domain adaptation is typically studied under two settings: unsupervised, where no target labels are available, and semi-supervised, where a small number of labeled target samples are accessible. EEG signals provide a classic example of domain shift: changes between calibration (training) and feedback (testing) sessions often degrade model performance. This has motivated adaptation methods such as covariate shift correction [4], which assume only that the marginal distribution $\mathbb{P}(X)$ changes, while $\mathbb{Q}(Y|X)$ remains stable. Optimal Transport (OT) provides a principled framework for domain adaptation by aligning distributions through the minimization of the transport cost, typically measured by the Wasserstein distance. In OT-based domain adaptation (OT-DA), the goal is to learn a push-forward map $\varphi:X_{S}\longmapsto X_{T}$ such that the pushforward distribution $\varphi_{\sharp}\mathbb{P}_{S}$ matches the target distribution $\mathbb{P}_{T}$. This is commonly achieved through regularized OT, allowing models trained in the source domain to generalize to the target domain while respecting the geometry of the data [5].
In Electroencephalogram (EEG)-based motor imagery classification, signals are commonly represented as spatial covariance matrices averaged over trials [6, 7]. Equipped with a Riemannian metric such as the Log-Euclidean metric, these covariance matrices lie on the manifold of SPD matrices, forming a smooth Riemannian manifold with a well-defined geometric structure [8, 9]. This structure provides a principled foundation for developing learning algorithms on manifolds [10]. When adapted to this geometry, OT-DA becomes particularly effective. For instance, Yair et al. [11] pioneered OT-DA on SPD manifolds by employing a squared Euclidean cost $c(x,\bar{x})=d_{\ell_{2}}(x,\bar{x})^{2}$, and derived a closed-form solution based on Brenier’s polar factorization. However, their method is limited by both theoretical soundness and practical performance:
  1. i.
    Euclidean cost fails to reflect the intrinsic Riemannian geometry of SPD manifolds.
  2. ii.
    The method presumes a stable $\mathbb{Q}(Y|X)$, overlooking label shift. Additionally, label scarcity in test phases and frequency-band variability pose further modeling challenges.
These limitations and challenges motivate the need for a more principled OT-DA framework that respects SPD geometry and accounts for both marginal and conditional distribution mismatches.
In this study, we focus on a domain adaptation scenario where the feature and label spaces are consistent ($X_{S}=X_{T}$, $Y_{S}=Y_{T}$), but both the marginal and conditional distributions differ. This setting calls for joint distribution adaptation, which simultaneously aligns $\mathbb{P}(X)$ and $\mathbb{Q}(Y|X)$, even when target labels are unavailable [12]. To model $X_{S}=X_{T}$ on the SPD manifold $\mathcal{S}_{++}$, we use the Log-Euclidean metric ($g^{LEM}$[8], which maps SPD matrices to a Euclidean tangent space at the identity matrix via matrix logarithms. This enables standard operations in the tangent space while preserving manifold structure. Notably, parallel transport under $g^{LEM}$ becomes identity, simplifying source-target alignment within a shared tangent space.
We propose Deep Optimal Transport (DOT), a geometric deep learning framework for joint domain adaptation on SPD manifolds that leverages a deep encoder to implicitly learn a nonlinear mapping on SPD manifolds as a geometric transformation. Specifically, the encoder-based paradigm projects source and target domains into a shared SPD subspace $Z\subset\mathcal{S}_{++}$ via $\varphi:X_{S},X_{T}\longmapsto Z$, aligning both marginal and conditional distributions through a composite optimal transport loss. The alignment is driven by minimizing the squared Log-Euclidean distance as the OT cost function, i.e., $c(x,\bar{x})=d_{g^{LEM}}^{2}(x,\bar{x})$, under the principle that domain discrepancies can be quantified and bridged through OT-based distribution matching. This enables $\mathbb{P}_{S}(\varphi(X_{S}))\approx\mathbb{P}_{T}(\varphi(X_{T}))$ and $\mathbb{Q}_{S}(Y_{S}|\varphi(X_{S}))\approx\mathbb{Q}_{T}(Y_{T}|\varphi(X_{T}))$ without assuming Gaussianity or linearity. We evaluate DOT on cross-session motor imagery classification, a challenging task where neuronal activity shifts over time, causing significant EEG signal variability across sessions. Our experiments span unsupervised and semi-supervised scenarios: the model is calibrated on one session and tested on another, demonstrating robustness to nonstationary brain dynamics. Unlike linear Wasserstein barycenter methods, DOT’s deep encoder captures nonlinear SPD embeddings, and the composite OT loss jointly addresses both domain-level and class-level distribution shifts.
@@ 1 Introduction
The remainder of this paper is organized as follows: Section 2 introduces basic concepts, including Log-Euclidean geometry, discrete OT, joint distribution alignment, and SPD matrix-based classifiers for motor imagery tasks. Section 3 details the proposed OT-DA framework on SPD manifolds and presents the architecture of DOT. Section 4 applies DOT to EEG-based motor imagery classification across diverse datasets. Section 5 analyzes results from synthetic and real-world experiments. A comparative summary of existing frameworks is provided in Table 1, highlighting key methodological distinctions.
Table 1: A comparison of various OT-DA Frameworks. $X_{S}$ and $X_{T}$ are feature spaces of the source and target domains, respectively. $\mathbb{P}$ and $\mathbb{Q}$ represent the marginal probability distribution and conditional probability distribution, respectively.
Framework Cost function Space Transformation Domain Adaptation Scenario
OT-DA[5] $c(x,\bar{x})=d_{\ell_{2}}^{2}(x,\bar{x})$ Euclidean Affine Function $X_{S}\neq X_{T}$.
JDOT-DA[13, 14] $c(x,\bar{x})=d_{\ell_{2}}^{2}(x,\bar{x})$ Euclidean Affine Function i). $X_{S}\neq X_{T}$;
$ii).\mathbb{Q}_{S}(Y_{S}|X_{S})\neq\mathbb{Q}_{T}(Y_{T}|X_{T})$.
OT-DA on $\mathcal{S}_{++}$[11] $c(x,\bar{x})=d_{\ell_{2}}^{2}(x,\bar{x})$ ($\mathcal{S}_{++}$, $g^{AIRM}$) BiMap Function $X_{S}\neq X_{T}$.
DOT (Proposed) $c(x,\bar{x})=d_{g^{LEM}}^{2}(x,\bar{x})$ ($\mathcal{S}_{++}$, $g^{LEM}$) Neural Network i). $X_{S}=X_{T}$;
ii). $\mathbb{P}_{S}(X_{S})\neq\mathbb{P}_{T}(X_{T});$
iii). $\mathbb{Q}_{S}(Y_{S}|X_{S})\neq\mathbb{Q}_{T}(Y_{T}|X_{T})$.
2 PRELIMINARY
@@
A matrix $S\in\mathbb{R}^{n\times n}$ is symmetric and positive-definite if it is symmetric and all its eigenvalues are positive, i.e., $S=S^{T}$, and $x^{T}Sx>0$, for all $x\in\mathbb{R}^{n}\setminus\{{\mathbf{0}}\}$. The Frobenius inner product and Frobenius norm on $m\times n$ matrices $A$ and $B$ are defined as $\langle\,A,B\,\rangle_{\mathcal{F}}:=\operatorname{Tr}{(A^{\top}\cdot B)}$ and $||\,A\,||_{\mathcal{F}}^{2}:=\langle\,A,A\,\rangle_{\mathcal{F}}$ respectively.
2.1 Riemannian Metric for SPD Manifolds
The Log-Euclidean metric ($g^{LEM}$), a widely adopted Riemannian metric on SPD manifolds, induces a vector space structure on the space of SPD matrices, enabling Euclidean-like operations in the logarithmic domain while preserving geometric properties [8, 9]. Formally, a commutative Lie group structure, known as the tensor Lie group, is endowed on $\mathcal{S}_{++}$ using the logarithmic multiplication $\odot$ defined as follows,
$P_{1}\odot P_{2}:=\exp\big{(}\log P_{1}+\log P_{2}\big{)},$
where $P_{1},P_{2}\in\mathcal{S}_{++}$, and $\exp$ and $\log$ are matrix exponential and logarithm respectively. The bi-invariant metric exists on the tensor Lie group, which yields the scalar product as follows,
$g^{LEM}_{P}(v,w):=\langle\,D_{P}\log v,D_{P}\log w\,\rangle_{\mathcal{F}},$
@@ 2.1 Riemannian Metric for SPD Manifolds
Another key Riemannian metric for SPD manifolds is the affine-invariant Riemannian metric ($g^{AIRM}$[15], which we compare with the Log-Euclidean metric. Formally, for any $P\in\mathcal{S}_{++}$, the $g^{AIRM}$ at $P$ is defined as:
$g^{AIRM}_{P}(v,w):=\langle\,P^{-\frac{1}{2}}\,v\,P^{-\frac{1}{2}},P^{-\frac{1}{2}}\,w\,P^{-\frac{1}{2}}\,\rangle_{\mathcal{F}},$
where tangent vectors $v,w\in\mathcal{T}_{P}\mathcal{S}_{++}$. The manifold ($\mathcal{S}_{++}$, $g^{AIRM}$) preserves critical geometric properties, such as affine invariance, scale invariance, and inversion invariance [16]. These properties make this metric theoretically appealing for covariance-based learning tasks.
2.2 Discrete Optimal Transport
@@
Discrete optimal transport aims to find the minimum-cost transportation plan between source and target distributions that are accessible only through finite sample observations [17]. Formally, given two measurable spaces $X$ and $\bar{X}$, the empirical distributions of measures $\mu\in X$ and $\nu\in\bar{X}$ are expressed as $\mu:=\sum_{i=1}^{N}p_{i}\delta_{x_{i}}$ and $\nu:=\sum_{j=1}^{M}q_{i}\delta_{\bar{x}_{i}}$, where the discrete variable $X$ takes values $\{x_{i}\}_{i=1}^{N}\in X$ with weights satisfying $\sum_{i=1}^{N}p_{i}=1$, while $\bar{X}$ assumes values $\{\bar{x}_{j}\}_{j=1}^{M}\in\bar{X}$ with weights $\sum_{j=1}^{M}q_{j}=1$. The Dirac measure $\delta_{x}$ represents a unite point mass concentrated at the point $x\in X$ or $\bar{X}$. The Kantorovich formulation of the discrete optimal transport problem is defined as
$\min_{\gamma\in\prod(\mu,\nu)}\big{\langle}\gamma,c(x,\bar{x})\big{\rangle}_{\mathcal{F}},$
where the transportation plan is constrained to the set $\gamma\in\prod(\mu,\nu):=\{\gamma\in\mathbb{R}_{\geq 0}^{|X|\times|\bar{X}|}\big{|}\gamma\mathbbm{1}_{|X|}=\mu,\text{ and }\gamma^{\top}\mathbbm{1}_{|\bar{X}|}=\nu\}$, and $\mathbbm{1}_{d}$ denotes a $d-$dimensional all-one vector.
2.3 Joint Distribution Optimal Transport
Assuming that the transport between source and target distributions can be estimated via optimal transport, Joint Distribution Optimal Transport (JDOT) [13] extends the OT-DA framework by jointly optimizing both the optimal coupling $\gamma$ and the prediction function $f$. This is formulated as follows,
$\min_{\gamma\in\prod(\mu,\nu)}\big{\langle}\gamma,c(x,y;\,\bar{x},f(\bar{x}))\big{\rangle}_{\mathcal{F}},$